Find a formula an for the nth term of the arithmetic sequence whose first term is a1=−3 such that an+1−an=4 for n≥1.


1
Expert's answer
2021-06-04T04:16:43-0400

Taking into account that "a_{n+1}=a_n+4," we conclude that the common difference of an arithmetic sequence is equal to 4. Therefore, the formula for the "n"th term is the following:

"a_n=4(n-1)+a_1=4n-4-3=4n-7."